# AI Foundation Models in Remote Sensing: A Survey

### v1

@@ III-A Self-Supervised Learning (SSL) Training Strategies for Remote Sensing FMs
Figure 3: General pipeline of SSL[35]. Diverse datasets images and pretext task images are acquired from ImageNet[13], BigEarthNet[60], and MillionAID[44]. Finetune dataset includes images from DIOR[38].
@@
 [^2] Object detection and Segmentation © MAXAR 2024, provided through the NextView License Agreement.
In remote sensing applications, two specific SSL methods used commonly include contrastive learning and predictive coding.
III-A1 Contrastive Learning.
Contrastive learning aims to learn representations by comparing different augmented views of the same data point. The main idea is to bring similar (positive) pairs closer and push dissimilar (negative) pairs apart in the feature space. This method relies heavily on data augmentation to create multiple views of the same image.
III-A2 Predictive Coding.
Predictive coding is another SSL technique where the model learns to predict missing parts of the input data from the observed parts. This method is beneficial for capturing spatial and temporal dependencies within the data. Common predictive coding methods include autoencoders (AE) and masked autoencoders (MAE).
Common SSL methods include SimCLR, MoCo (Momentum Contrast), BYOL (Bootstrap Your Own Latent), and DINO (Self-Distillation with No Labels). Each of these methods has a distinctive approach to creating positive and negative pairs and updating model parameters. These methodologies have proven effective in tasks such as scene classification, semantic segmentation, and object detection in remote sensing. For example, SSL can help models learn to classify land cover types, identify objects like buildings and vehicles, and segment different regions within satellite images, even when labeled data is limited.
III-B Backbones
In deep learning, backbones are essential neural network architectures that serve as feature extractors. They form the foundational layers of a model, processing input data to generate rich, hierarchical feature representations. These representations can then be utilized by subsequent model components to perform various tasks such as classification, detection, and segmentation. By leveraging powerful backbones, models can efficiently handle complex data and improve their performance across different applications. In the following section, we
III-B1 Backbone Type I: Transformers and Vision Transformers (ViT)
Transformers, originally designed for natural language processing, have revolutionized the field of computer vision with their ability to model long-range dependencies through self-attention mechanisms. Vision Transformers (ViT) [19] adapt the transformer architecture proposed by Vaswani et al. (2017)[67] to handle image data by treating image patches as sequences of tokens. This adaptation is particularly useful in remote sensing, where images can be large and contain complex spatial structures. Figure 3 illustrates the basic structure of ViT for RS segmentation tasks.
The key components of ViT include patch embedding, positional encoding, the transformer encoder, and the classification head [19]. Patch embedding involves dividing an image into fixed-size patches and linearly embedding each patch into a vector. Positional encoding adds positional information to the patch embeddings to retain the spatial structure of the image. The transformer encoder consists of multiple layers of multi-head self-attention and feed-forward neural networks, which process the sequence of embedded patches to capture global dependencies. Finally, the classification head is a fully connected layer that processes the final sequence representation for downstream tasks such as image classification. The self-attention mechanism in transformers allows each token to attend to all other tokens, providing a powerful way to capture global context. The attention weights are computed as:
@@ III-B1 Backbone Type I: Transformers and Vision Transformers (ViT)
where $Q$ (query), $K$ (key), and $V$ (value) are the input matrices, and $d_{k}$ is the dimension of the key vectors [67]. ViTs have shown great promise in remote sensing tasks such as land cover classification, urban area identification, and vegetation analysis, leveraging their ability to capture both local and global patterns.
Figure 4: The Vision transformer architecture.
 [^3] RGB Image and Segmentation © MAXAR 2024, provided through the NextView License Agreement.
III-B2 Backbone Type II: Convolutional Neural Network (CNN)
Convolutional Neural Networks (CNN), such as Residual Neural Networks (ResNet), address the degradation problem in deep neural networks by introducing residual connections, which allow gradients to bypass certain layers, facilitating the training of very deep networks. This capability is particularly beneficial in remote sensing, where deep models are often required to capture the intricate details and variations in satellite images.
ResNet, as an example, is characterized by their residual blocks, which include shortcut connections that bypass one or more layers. The residual block can be described by the following equation:
$\mathbf{y}=\mathcal{F}(\mathbf{x},\{W_{i}\})+\mathbf{x}$
where $\mathbf{y}$ is the output, $\mathcal{F}$ represents the residual mapping to be learned, $\mathbf{x}$ is the input, and $\{W_{i}\}$ are the weights of the layers in the block[26]. Depending on the dimensions, shortcuts can be identity mappings (if input and output dimensions match) or convolutional layers (if dimensions differ).
ResNet has various architectures like ResNet-50, ResNet-101, and ResNet-152, with the number indicating the total layers. These networks have shown remarkable performance in various vision tasks due to their ability to train deeper networks without degradation. In remote sensing, ResNets are widely used for image classification, object detection, and change detection tasks. For example, ResNet-based models can classify different land cover types, detect objects like buildings and vehicles, and monitor changes in the landscape over time by comparing temporal sequences of satellite images.

### v2

@@ IV Pretraining Methods
Figure 3: General pipeline of SSL[51]. Diverse datasets images and pretext task images are acquired from ImageNet[18], BigEarthNet[85], and MillionAID[64]. Finetune dataset includes images from DIOR[55].  [^4] Object detection and Segmentation © MAXAR 2024, provided through the NextView License Agreement.
Pretraining serves as a critical step in developing foundation models (FM), enabling them to learn transferable and generalized representations from large-scale datasets. This process leverage self-supervised or supervised learning methods to extract domain-agnostic features that can be adapted to various downstream tasks. In this section, we explore the key pretraining methods utilized commonly in foundation models for remote sensing, explaining the mechanism of these methods and their roles in enhancing model performance and addressing challenges in this field.
IV-A Self-Supervised Learning
Self-supervised learning has emerged as a cornerstone of pre-training foundation models, offering a paradigm where models learn representations by predicting parts of the input data from other parts. This approach reduces reliance on expensive and time-consuming labeled datasets, making it particularly advantageous in fields like remote sensing, where labeled data is often scarce or challenging to obtain.
SSL allows models to exploit vast amounts of unlabeled data, learning rich, generalizable representations that transfer well to downstream tasks such as scene classification, semantic segmentation, object detection, and change detection. By uncovering underlying data structures and patterns, SSL not only enhances model robustness but also improves adaptability across diverse domains and resolutions of remote sensing imagery [104]. Figure 4 illustrates the general pipeline of self-supervised learning.
Two SSL methods commonly used in vision foundation models for remote sensing are contrastive learning and predictive coding, each offering unique mechanisms to harness information from unlabeled data.
IV-A1 Predictive Coding
Predictive coding leverages a generative approach, where the model learns to predict missing or occluded parts of an image based on visible portions. This strategy helps capture spatial and contextual relationships in remote sensing imagery, which often contains diverse textures, complex scenes, and varying resolutions.
In remote sensing, predictive coding can be applied to tasks such as gap filling in satellite imagery, where the model learns to infer missing data caused by sensor limitations or occlusions like cloud cover. Popular implementations of predictive coding frameworks include autoencoder-based architectures, masked image modeling techniques like those used in MAE (Masked Autoencoders) [34], and autoregressive models. These methods are particularly effective in learning fine-grained details critical for high-resolution imagery and specialized tasks.
IV-A2 Contrastive Learning
Contrastive learning is another powerful SSL technique that focuses on distinguishing between similar and dissimilar samples in the data. The key idea is to bring representations of similar (positive) samples closer together while pushing apart those of dissimilar (negative) samples. This encourages the model to learn discriminative and invariant features that are crucial for remote sensing tasks.
Contrastive learning frameworks such as SimCLR [13], MoCo [35], DINO [9], and BYOL [29] have shown promise in remote sensing applications. They use augmentations like random cropping, rotations, or spectral band dropping to generate positive pairs, enabling the model to learn robust representations invariant to these transformations. For instance, in multispectral or hyperspectral imagery, contrastive learning can help models capture spectral signatures across varying conditions, improving performance in tasks like crop classification or land cover mapping [104].
Contrastive learning is especially relevant in remote sensing when labeled datasets are highly imbalanced, as it enables models to learn from underrepresented classes or regions without explicit labels.
By combining approaches like predictive coding and contrastive learning, self-supervised learning has significantly advanced the development of vision foundation models in remote sensing. These methods allow models to leverage vast, unlabeled datasets while maintaining adaptability across diverse spatial resolutions, spectral bands, and application scenarios. On the other hand, it is important to note that there are many other SSL methods that can be employed to such tasks. Other innovative methods, such as teacher-student self-distillation frameworks, have also demonstrated potential in remote sensing applications. For example, CMID [70] achieves promising performance by combining contrastive learning and masked image modeling in a teacher-student self-distillation framework. This structure enables it to capture both global and local features, making it effective for diverse remote sensing tasks.. The diversity of SSL techniques highlights the versatility and evolving nature of self-supervised learning, underscoring its critical role in unlocking the full potential of remote sensing imagery.
IV-B Supervised Pretraining
Supervised pretraining is a fundamental approach in deep learning, where models are trained using labeled datasets to minimize prediction errors for specific tasks, such as image classification. This method allows models to learn direct mappings between input features and target labels, fostering the development of detailed, task-specific representations. For instance, models like ResNet [36] and VGGNet (Visual Geometry Group Network) [81] trained on large-scale datasets such as ImageNet [18] have demonstrated how supervised pretraining can capture robust feature hierarchies that are highly transferable to related tasks, including semantic segmentation or object detection.
In remote sensing, supervised pretraining has shown promise for tasks such as land cover classification and object detection using high-resolution satellite imagery [97]. However, the dependency on large-scale labeled datasets presents a major limitation. Creating labeled datasets for remote sensing tasks, particularly when involving multispectral or hyperspectral data, is resource-intensive and often requires domain expertise for annotation. For example, labeling pixel-level data for land cover classification or delineating objects in complex urban environments can be prohibitively time-consuming. Furthermore, labeled data in remote sensing is often domain-specific, limiting the generalizability of models trained on one dataset to other applications or regions [129].
These challenges highlight the need for innovative strategies to address the reliance on labeled data. Such limitations have motivated the development of alternative approaches, including self-supervised pretraining methods, which leverage the abundance of unlabeled data to learn general-purpose representations without manual annotation.
V Image Analysis Methods
V-A Image Perception at Different Levels
Foundation models in remote sensing enable image analysis at three primary levels: image-level, region-level, and pixel-level. These levels address varying spatial, contextual, and application-specific needs, providing the foundation for a wide range of tasks such as environmental monitoring, urban planning, disaster response, and more. The following subsections outline the distinct objectives and applications at each level. A detailed summary of the models evaluated for these tasks is provided in table LABEL:tab:cv-task. The following subsections outline the distinct objectives and applications at each level.
V-A1 Image-Level
Image-level analysis focuses on classification tasks, categorizing entire images or large image segments into predefined classes, such as urban, forest, water bodies, or agricultural areas. This approach provides broad, high-level insights into geographic regions and is instrumental in large-scale applications like land use mapping, land cover classification, and resource management. By classifying entire scenes, this level of analysis enables efficient monitoring of extensive areas, supporting decision-making in environmental management and policy planning.
V-A2 Region-Level
Region-level analysis identifies and localizes specific objects within an image, such as buildings, vehicles, ships, or other structures. Unlike image-level analysis, which provides holistic classifications, region-level tasks focus on object detection which is to detect individual entities and their spatial locations. This analysis is critical for targeted applications like urban planning, where the detection of infrastructure is essential, as well as disaster response and security, where identifying damaged buildings or vulnerable areas can significantly aid in timely interventions.
TABLE II: Overview of recent foundation models in remote sensing, categorized by architecture, model name, pre-training dataset, resolution, geographic coverage, image analysis lebels, visual encoder, pre-training methods, and the number of parameters. Abbreviations for pre-training methods as specified in the original work: SSL refers to Self-Supervised Learning, CL stands for Contrastive Learning, MIM is Masked Image Modeling, MAE is Masked Autoencoders, and FD-MIM is Feature-Distilled Masked Image Modeling.
Model Name Architecture Pre-training Dataset Resolution (m) Geographic Coverage
Image Analysis
Levels
Pretrain methods # of Params
CMC-RSSR [84] ResNet-50
NWPU-DOTA[114],
BigEarthNet [85],
ImageNet [18]
0.2 to 60 Global
Image-level
Contrastive
Multiview Coding
23M
SeCo [66] ResNet-50 Sentinel-2 Imagery 10, 20, 60
200k Locations
Worldwide
Image-level,
Spacial-temporal
CL 23.5M
GeoKR [56] ResNet-50 Levir-KR [56] 0.8 to 16 Global
Image-level, Pixel-level,
Region-level
Geographical Knowledge
Supervision
23.5M/138M
MATTER [2] ResNet-34 Sentinel-2 Imagery -
Rural and Remote Regions
with Little Changes
Image-level,
Pixel-level
SSL 21.3M
GASSL [6] ResNet-50 fMoW [15], GeoImageNet [18] - 7 Continents
Image-level, Pixel-level
Region-level
CL 23.5M
RSP [96] ViTAEv2-S MillionAID [64, 63] 0.5 to 153 Global
Image-level, Pixel-level,
Region-level, Spacial-temporal
Supervised Learning 24.8M/23.5M/29M
DINO-MM [105] ViT-S/8 BigEarthNet-MM [86] 10 Global Image-level SSL 22M
Scheibenreif, et al. [79] Swin Transformer SEN12MS [80] 10 Global
Image-level,
Pixel-level
CL -
RingMo [87] ViT/Swin Transformer 2 million RS images 0.3 to 30 6 Continents
Image-level, Pixel-level,
Region-level, Spacial-temporal
MIM -
GeCO [57] ResNet-50 Levir-KR [56] 0.8 to 16 Global
Image-level, Pixel-level,
Region-level
SSL 23.5M
RS-BYOL [45] BYOL Sen12MS [80] 10 to 20 Global
Image-level,
Pixel-level
SSL 23.5M
CSPT [124] ViT-B ImageNet-1K [18] - Global
Image-level,
Region-level
SSL 86M
RVSA [100] ViT MillionAID [64, 63] 0.5 to 153 Global
Image-level, Pixel-level,
Region-level
MAE 100M
SatMAE [16]
MAE-based
Framework
fMoW Sentinel-2 [15] 10, 20, 60 Global
Image-level,
Pixel-level
MAE 307M
TOV [89] TOV
TOV-NI,
TOV-RS
- Global
Image-level, Pixel-level,
Region-level
SSL -
CMID [70]
Teacher-student
Self-distillation
MillionAID [64, 63] Varied Global
Image-level, Pixel-level,
Region-level, Spacial-temporal
SSL 25.6M/87.8M
CACo [67] ResNet-18/50 Sentinel-2 Imagery 10 Global
Image-level, Pixel-level,
Spacial-temporal
SSL 11.7M/23.5M
IaI-SimCLR [77] ResNet-18 SEN12MS - Global Image-level CL 11.7M
SSL4EO-L [83] ResNet/ViT
ImageNet [18],
MoCo [35], SimCLR [13]
30 Global Pixel-level SSL 11.7M/23.5M/86M
GFM [69] Teacher-Student GeoPile [69] - Global
Image-level,
Pixel-level
Continual Pretraining -
SatlasPretrain [7] SatlasNet GeoPile [69] 1, 10 Global
Image-level,
Pixel-level
Multi-task Learning 88M
RingMo-Sense [119] Multi-Branch
RS Spatiotemporal
Dataset
- Global Pixel-level SSL -
Scale-MAE [78] ViT-Large FMoW [15] - Global
Image-level,
Pixel-level
MAE 322.9M
RingMo-lite [109] CNN-Transformer AID [115] 0.3 to 30 Global
Image-level,
Pixel-level,
Region-level,
Spacial-temporal
FD-MIM
60%
less than RingMo
DeCUR [102] Multimodel SSL
SSL4EO-S12 [107],
RGB-DEM/depth
Varied Global
Image-level,
Pixel-level
SSL 23.5M
Feng et al. [27] MSFE+MMFH Multi-modal Dataset Varied Global
Image-level,
Pixel-level,
Region-level,
Spacial-temporal
SSL -
FG-MAE [108] ViT SSL4EO-S12 [107] 10 Global
Image-level,
Pixel-level
MAE -
Prithvi [46] ViT Harmonized Landsat Sentinel 2 30 Contiguous U.S. Pixel-level MAE 100M
CROMA [28] Multimodal Encoder SSL4EO [107] 10
Areas Surrounding
Human Settlements
Image-level,
Pixel-level
CL, MAE 86M
USat [44] ViT Satlas [7] Varied Global Pixel-level MAE -
Cross-Scale MAE [88] ViT-B fMoW [15] - Global
Image-level,
Pixel-level
MAE 86M
U-BARN [26] Unet+Transformer Sentinel-2 Imagery Varied France
Image-level,
Pixel-level
SSL -
EarthPT [82] Transformer Sentinel-2 Imagery 10 UK Image-level Autoregressive SSL 700M
GeRSP [42]
Teacher-Student
Network
ImageNet [18],
MillionAID [64, 63]
0.5 to 153 Global
Image-level,
Pixel-level,
Region-level
SSL, SL -
SwiMDiff [91] Dual-Branch Sen12MS [80] Varied Global
Image-level,
Spacial-temporal
SSL 11.7M
SMLFR [22] Generative ConvNet GeoSense [23] 0.05 to 150 Multiple Continents
Pixel-level,
Region-level
SSL 88M/197M
SpectralGPT [40] 3D GPT Sentinel-2 Imagery Varied Global
Image-level,
Pixel-level,
Spacial-temporal
MAE 100M/300M/600M
Presto [92]
MAE-based
Framework
Presto-21.5M [92] 10 Global Crop Type Segmentation MAE 402K
SatMAE++ [73] SatMAE fMoW [15] Varied Global Image-level
Multi-Scale Pre-training
SAR-JEPA [58]
Joint-Embedding
Predictive Architecture
100K SAR Images Varied Global Image-level SSL -
FoMo-Bench [8] ViT Multiple Varied Global
Image-level,
Pixel-level, Region-level
MAE 101M/110M
SkySense [32]
Factorized Multi-Modal
Spatiotemporal Encoder
Multiple Varied Global
Image-level,
Pixel-level,
Region-level,
Spacial-temporal
CL 2.06B
UPetu [24] Multi-Modules GeoSense [23] - Global
Image-level,
Pixel-level, Spacial-temporal
SSL 0.65M
msGFM [33] Swin Transformer GeoPile-2 [69] 0.1 to 153 Global
Image-level, Pixel-level
MIM 89M
DINO-MC [111] DINO SeCo-100K [66] 10 to 60 Global
Image-level, Spacial-temporal
SSL -
OFA-Net [118] OFA-Net Multi-modal Dataset Varied Global
Image-level, Pixel-level
MIM -
MTP [99]
Shared Encoder
Task-Specific Decoders
SAMRS [98] Varied Global
Image-level, Pixel-level,
Region-level, Spacial-temporal
Multi-Task
Pretraining
over 300M
BFM [11] ViT MillionAID [64, 63] 0.5 to 153 Global
Pixel-level,
Region-level
MAE
86M/605.26M/
1.36B/2.42B
MMEarth [72] MP-MAE
Multi-modal,
Geospatial Data
- Global
Image-level, Pixel-level
MP-MAE 3.7M to 650M
CtxMIM [123] ViT WorldView-3 Imagery Varied Asia
Image-level,
Pixel-level,
Region-level
MIM 88M
SARATR-X [54] HiViT SAR Datasets 0.1 to 3 Global
Image-level, Region-level
MIM 66M
SoftCon [106]
Siamese Network
with ResNet
and ViT Backbones
SSL4EO-S12-ML [107] - Global
Image-level,
Pixel-level,
Spacial-temporal
Multi-label Soft
Contrastive Learning
23M, 23M, 86M
LeMeViT [49] Hierarchical ViT MillionAID [64, 63] - -
Image-level, Pixel-level,
Region-level , Spacial-temporal
Dual Cross-Attention
with Learnable Meta Token
Adaptation
8.33M to 52.61M
S2MAE [59]
3D Transformer-based
MAE
fMoW-Sentinel [15],
BigEarthNet [85]
- Global
Image-level, Spacial-temporal
3D MAE -
RS-DFM [110]
Multi-platform
Inference Framework
AirCo-MultiTasks [110] - -
3D Region-level,
Pixel-level
Generalized Feature
Mapping with Relative
Depth Estimation
A2-MAE [122] ViT-Large
STSSD
(Spatial-Temporal-Spectral
Structured Dataset)
0.8 - 30m Global
Image-level,
Pixel-level,
Spacial-temporal
Anchor-aware
Masking Strategy
and Geographic
Encoding Module
304M
HyperSIGMA [95] ViT-based HyperGlobal-450K [95] 30m Global
Image-level, Region-level,
Anomaly Detection,
Spacial-temporal
MAE over 1B
DOFA [117] Dynamic OFA Multiple 1 to 30 Global
Image-level,
Pixel-level
MIM  [^1] M/337M
V-A3 Pixel-Level
Pixel-level analysis offers the most granular form of image perception, assigning a label to every pixel within an image. This includes tasks such as semantic segmentation, where each pixel is classified into categories like vegetation, water, or buildings; it also include change detection, which identifies temporal differences between images captured at different times. Pixel-level analysis is indispensable for creating highly detailed maps used in applications like precision agriculture, deforestation tracking, and disaster management. The ability to analyze fine-grained details enables more accurate assessments and actionable insights for these critical areas.
@@ V-A3 Pixel-Level
Figure 4: The Vision transformer architecture. [^6] RGB Image and Segmentation © MAXAR 2024, provided through the NextView License Agreement.
V-B Backbone
V-B1 Convolutional Neural Networks (CNNs)
Convolutional Neural Networks [74] are a fundamental architecture in deep learning, designed to extract hierarchical spatial features from images through the use of convolutional layers. Each convolutional layer applies filters to the input data, detecting patterns like edges, textures, and shapes at different levels of abstraction. This makes CNNs well-suited for handling complex visual tasks in remote sensing, such as image classification, segmentation, and object detection.
Residual Neural Networks (ResNet) [36], a type of Convolutional Neural Network (CNN) , address the degradation problem in deep neural networks by introducing residual connections, which allow gradients to bypass certain layers, facilitating the training of very deep networks. This capability is particularly beneficial in remote sensing, where deep models are often required to capture the intricate details and variations in satellite images. ResNet, as an example, is characterized by their residual blocks, which include shortcut connections that bypass one or more layers. The residual block can be described by the following equation:
$\mathbf{y}=\mathcal{F}(\mathbf{x},\{W_{i}\})+\mathbf{x}$
where $\mathbf{y}$ is the output, $\mathcal{F}$ represents the residual mapping to be learned, $\mathbf{x}$ is the input, and $\{W_{i}\}$ are the layer weights [37].
ResNet has various architectures like ResNet-50, ResNet-101, and ResNet-152, with the number indicating the total layers. These networks have shown remarkable performance in various vision tasks due to their ability to train deeper networks without degradation. In remote sensing, ResNets are widely used for image classification, object detection and change detection tasks [31] . For example, ResNet-based models can classify different land cover types [128, 125], detect objects like buildings and vehicles [31], and monitor changes [128, 75] in the landscape over time by comparing temporal sequences of satellite images.
V-B2 Transformers and Vision Transformers (ViTs)
Transformers, adapted for computer vision (CV) as Vision Transformers (ViT), model long-range dependencies through self-attention, making them effective for complex geospatial data. Figure 6 illustrates the architecture of ViT. ViTs treat images as sequences of patches, capturing global and local patterns, which is useful for segmentation and change detection. The self-attention mechanism computes:
@@ V-B2 Transformers and Vision Transformers (ViTs)
where $Q$ (query), $K$ (key), and $V$ (value) are the input matrices, and $d_{k}$ is the dimension of the key vectors [93].